# Laser-driven Ultrafast Dynamics of a Fractional Quantum Hall System

### v1

@@ Abstract
Where $X_{j}=\frac{2\pi j}{L_{y}}\ell_{B}^{2}$ with $\ell_{B}=\sqrt{\hbar/eB}$ being the magnetic length. ${\cal N}_{n}=L_{y}\sqrt{\pi}2^{n}n!\ell_{B}$ is the normalization constant and $H_{n}$ is Hermite polynomial of order $n$. For each Landau level $n$, index $j$ runs from $10,1,...,N_{\phi}-$ 1, where $N_{\phi}=L_{x}L_{y}/(2\pi\ell_{B}^{2})$ is the number of flux quanta (Landau orbitals) passing through the system. Within this Landau orbital basis, we are able to write the system Hamiltonian as $\hat{H}=\hat{H}_{sp}+\hat{H}_{int}$, with
@@
Figure 1: (a) Torus with lengths $L_{x}$ and $L_{y}$. Magnetic field is pointing along $\hat{z}$. (b) A schematic of a two-level FQH system. Initially, the system is in fractional $\nu=1/3$ state occupying the LLL ($n=0$). This state is then laser-driven creating excitations to the first Landau level ($n=1$). The energy difference between the two levels is $\omega_{c}$.
In the absence of the pulse field and for large positive $\omega_{c}$, our fermions are in the ground state of the system, containing $N$ electrons, such that the filling factor is $\nu{=}N/N_{\phi}{=}1/3$. Also the system Hamiltonian commutates with the center of mass-momentum operator $\hat{K}=\sum_{j,n}jc_{n,j}^{\dagger}c_{n,j}$ (in Landau gauge) and momentum transfer by the laser pulse is considered to be negligible. Hence, our many-body dynamics is contained in the momentum eigen-sector of the initial state.
Figure 2: Two-electron virtual processes that do not conserve Landau level index.
To include the effects of the out-of-equilibrium physics under the influence of a laser-pulse, the Hamiltonian includes both scattering processes that conserve Landau level index $n_{1}+n_{2}=n_{3}+n_{4}$ and ones that break the level index i.e. $n_{1}+n_{2}\neq n_{3}+n_{4}$. The latter is schematically illustrated in Fig. 2. This is the consequence of an electron scattering from one level to another and is ignored in traditional and synthetic bi-layers Ghazaryan et al. (2017). We avoid large scale Hilbert space by truncating Hamiltonian to LL1 and by considering low-intensity radiation. This assumption is further validated by another resonant inelastic light experiment for FQH $\frac{1}{3}$ state showing most features related to excitations from LLL to LL1 Pinczuk et al. (1993). Coulumb’s interaction potential for torus boundary conditions is $V(\mathbf{r})=\sum_{s,t}e^{2}/\varepsilon|\mathbf{r}+sL_{x}\hat{x}+tL_{y}\hat{y}|$ where $\hat{x}$, $\hat{y}$ are the units vectors along the two axes of the torus Yoshioka et al. (1983). We obtain the matrix elements
@@
Where $C_{0,0}(\mathbf{q})=1$, $C_{1,1}(\mathbf{q})=1-\frac{|q|^{2}\ell_{B}^{2}}{2}$ and $C_{1,0}(\mathbf{q})=\frac{q_{y}+iq_{x}}{\sqrt{2}}\ell_{B}=\big{[}C_{0,1}(-\mathbf{q})\big{]}^{*}$. Prime over a delta function indicates that it is defined modulo $N_{\phi}$. Fractional $1/3$ state Laughlin (1983) is the ground state of the Hamiltonian $\sum_{j_{i}}V_{j_{1},j_{2},j_{3},j_{4}}^{0,0,0,0}c_{j_{1}}^{\dagger}c_{j_{2}}^{\dagger}c_{j_{3}}c_{j_{4}}$ and is adiabatically connected to period-three charge-density $|100,100,100,100,...\rangle$ Seidel et al. (2005); Trugman and Kivelson (1985) (for short-range potential $\nabla^{2}\delta(\mathbf{r})$). The physics of FQH systems are usually studied in the limit of large level-spacing and the Hamiltonian for single Landau level is diagonalized. This is due the assumption that Coulomb interaction is weak to scatter electron from one level to another in the large magnetic field limit which is not always true Papić and Balram (2022). We have first verified that the ground state of of the system Hamiltonian involving virtual process for various values of level spacing is indeed the fractional-$\frac{1}{3}$ quantum Hall state (See Supplemental Material (SM) sup for more details).
@@
$\hat{J}_{x/y}=\frac{\hbar(i/1)}{m\sqrt{2}}\sum_{j}\bigg{(}c_{1,j}^{\dagger}c_{0,j}\mp c_{0,j}^{\dagger}c_{1,j}\bigg{)}\;.$ (5)
The details of this quantity is given in the SM sup .
Figure 3: Color online. Dynamical observables for various laser frequencies $\omega_{0}$ (a) Number of excited electrons for various in-coming light frequencies in-set of (a) gives the form of $\Gamma_{\rm EM}(t)$ (b) Fidelity $|\langle\Psi(0)|\Psi(t)\rangle|$. (c) The y-component of current from Eq. (5). d) $x$-component of the current. The quantities are obtained for $N=6$ on torus with $L_{2}=10.63\ell_{B}$. The intensity is set to $eA_{0}/\sqrt{2}m^{*}\ell_{B}=0.2E_{c}$ and $t_{d}=10/E_{c}$. The units of time is the inverse of $E_{c}$. The usual values of $E_{c}$ is around $14\;{\rm meV}$ Pinczuk et al. (1993) which gives $47$ femto-seconds as the unit of time.
Figure 4: Spectrum of $d[\hat{J}_{y}(t)]/dt$ following the pulse with parameters $eA_{0}/\sqrt{2}m^{*}\ell_{B}=0.2E_{c}$ and $t_{d}=10/Ec$. The in-coming laser frequency is swept from $\omega_{0}=0.6-0.8\omega_{c}$. The behaviour of the current shows the excitation of a mode in addition to magnetoplasmon at frequency $w_{c}$ that is blue-shifted from $w_{c}$. The system size is set to $N=6$ and $L_{y}=\sqrt{2\pi 18}\ell_{B}$. Inset gives the spectrum for near-TT limit ($L_{y}=2\pi\ell_{B}$).
Results. We consider level-spacing $\omega_{c}<E_{c}$, i.e. the level spacing is less than Coulomb energy scale. The ratio of these energies $\kappa\equiv\frac{e^{2}}{\hbar\omega_{c}\epsilon\ell_{B}}$ is magnetic field dependent ranging from $2.6/\sqrt{B[T]}$ for Ga-As to $22.5/\sqrt{B[T]}$ for Al-As justifying the our approximation for fractional $\nu=1/3$ Hall system usually observed at $B\sim 10$ Tesla Sodemann and MacDonald (2013). In Fig. 3, we present observables like fidelity, mean-excitation number and spectrum of $d[\hat{J}(t)]/dt$ after the pulse for $\omega_{c}=E_{c}/2$ for $L_{y}=\sqrt{2\pi 18}\ell_{B}$. Fig. 3 (a-b) gives the excitations to LL1 and the fidelity. The current $\hat{J}_{x,y}$ in the Fig. 3 (c-d) remains oscillating with multiple periods after the pulse is finished. Fig. 4 gives the spectrum of current of ${\mathcal{J}}_{y}(\omega)={\mathcal{F}}_{t}\big{[}\frac{d}{dt}\hat{J}_{y}(t)\big{]}(\omega)$ for $L_{y}=\sqrt{2\pi 18}\ell_{B}$ (liquid limit) and $L_{y}=2\pi\ell_{B}$ (near-TT limit). There are two modes as seen from the spectrum for low-intensity short time pulses one at $\omega_{c}$ and another blue-shifted from $\omega_{c}$. The presence of these modes are robust against slight change in the incoming frequencies although their strength can change. The peak at the level-spacing (cyclotron-energy) $\omega_{c}$ is the plasmon equivalent of FQH state identified as $q\to 0$ limit of magnetoplasmon mode. This mode is due to single particle excitations from LLL to LL1 Pinczuk et al. (1993). In the experiments, another peak that is blue shifted from magnetoplasmon has also been observed Pinczuk et al. (1993). This mode was sometimes attributed to $q=0$ spin-flip inter-Landau-level excitation involving enhanced exchanged in spin-polarized state. We find blue-shifted peaks in our calculations are due to excitations of additional electron to $LL1$ although complete understanding is missing due to the presence of finite-size effects and the truncation to LL1.
@@
Figure 5: Color online. Structure factor ${\mathcal{S}}(q_{x}=2\pi/L_{x},\omega)$ for different system sizes. The torus dimensions is taken $L_{x}=L_{y}=\sqrt{6\pi N}\ell_{B}$. The laser-frequency is $\omega_{0}=0.7\omega_{c}$ and other laser parameters are set as in Fig. 3.
@@
The spectrum of structure factor ${\mathcal{S}}(\mathbf{q},\omega)$ is shown in Fig. 5. It shows the existence of low-energy modes for different system sizes. The peak at $\omega/\omega_{c}=1$ is magnetoplasmon mode discussed in the laser-field induced current spectrum. Significantly, there emerge new modes located at the energy around $0.2\omega_{c}$. We have also verified that in the low intensity regime of the laser pulse field, the magnetoplasmon mode and the new mode follow the linear and quadratic dependence of the field strength, respectively. To determine the nature of theses new low-energy modes in $S_{\mathbf{q}}(\omega)$, we evaluate the overlap $|\langle{\Psi(t)}|n\rangle|$ to analyze the weight of low-lying eigen-states $|n\rangle$ contained in the photo-induced excited state $|\Psi(t)\rangle$. The results are given in Table 1. We have found that low-energy excitations are dominated by quadrupole FQH graviton states. These states are identified by matrix-element $I_{n}=|\langle\Psi_{0}|\hat{O}^{(2)}|n\rangle|$, where $\hat{O}^{(2)}$ is the operator with quadrupole structure and $|\Psi_{0}\rangle$ is our initial ground state before the application of laser pulse. The specific form of $\hat{O}^{(2)}$ is given as Liou et al. (2019)
@@
Figure 6: Color online. (a) $I_{n}=|\langle\Psi_{0}|\hat{O}^{(2)}|n\rangle|$ for near-TT limit $L_{y}=2\pi\ell_{B}$. Where $|\Psi_{0}\rangle$ is the FQH-1/3 state and $|n\rangle$ is the $n$-th excited state of Hamiltonian in Eq. (3). The inset of (a) gives the transition matrix element between FQH-1/3 state and the eigen-states of LLL. (b) Same as (a) but for quantum-liquid limit ($L_{y}=10.63\ell_{B}$) for square torus. The inset of (b) gives the matrix element for the liquid-limit in LLL. The system size is $N=6$ particles and the level-spacing is set to $0.5E_{c}$. Grey vertical line on plots is at energy $0.15E_{c}$.
In Fig. 6, we present the values of transition matrix element $\langle n|\hat{O}^{(2)}|\Psi_{0}\rangle$ (un-normalized) between FQH 1/3 state and $|n\rangle$ is the $n$-th excited state for two cases i) When $|n\rangle$ is the eigen-state of the system Hamiltonian involving all scattering processes and ii) when $|n\rangle$ is the eigen-state of LLL Hamiltonian where all scattering processes are confined to LLL. This transition matrix element represents the system transition rate due to an oscillating metric tensor analogous to the passing of a gravitation wave and hence usually termed as graviton Yang (2016). This graviton mode occurs for both near-TT and liquid limit as shown in Fig.  6 (a) and (b). It is robust against the system size (See SM  sup for details). Moreover, as seen from Fig. 6, the energy of graviton states $E_{grav}$, when virtual processes are included, is slightly lower than the graviton states of a system when all interaction processes are confined to LLL. Furthermore, we follow Ref. Liou et al., 2019 to analyze the handedness of this graviton and find it has the structure of $(q_{x}-iq_{y})^{2}$, suggesting its chiral nature (See SM  sup for the detail of analysis.). Previously, it has also been proposed  Kirmani et al. (2022); Liu et al. (2018) that graviton can be excited via geometric quench Liu et al. (2021). A variant of such graviton mode called chiral graviton mode has been observed in experiments recently for fractional Hall states at filling fractional $\frac{1}{3}$ and $\frac{2}{5}$ fillings Liang et al. (2024). Chiral graviton modes in FQH system has been previously excited by tuning the laser-frequency near resonance conditions between valence and conduction band in 2D quantum wells. Our work has demonstrated that chiral graviton modes can also be exited when the a short-time laser is tuned near Landau level gap which is easily accessible experimentally. Furthermore, the excitation of graviton mode can be observed along with magneto-plasmon mode when all scattering processes are considered. We suggest to use a circularly polarized probe light to experimentally detect the chirality of the graviton mode.
@@
Table 1: The contribution of states in $|\Psi(t)\rangle$ after the pulse for $N=5$ and $N=6$. $\Delta E$ denotes the energy difference from the ground state. Overlap $|\langle n|\Psi(t)\rangle|$ gives the contribution of individual eigen-state $|n\rangle$ of the system Hamiltonian with energy $E_{n}$. $\omega_{c}=0.5E_{c}$.
@@
Size  [^1] The parameters of laser is set to $(eA_{0}/\sqrt{2}m^{*}\ell_{B},\omega_{0})=(0.2E_{c},0.7\omega_{c})$ with $t_{d}=10E_{c}$ and $L_{y}=\sqrt{2\pi N_{\phi}}\ell_{B}$. Similar results have been found for $N=7$ and are also observed for near-TT limit. $E_{n}$ ($E_{c}$) Overlap $\Delta E$ $(E_{c})$ Nature  [^2] Due to large Hilbert space size, we do not have eigen-state corresponding to Blue-shifted peak for $N=6$. Similarly, for $N=7$ we only have low-energy quadrupole states.
$N=6$1.38 0.84 0.0 G.S.
- -1.296 0.092 0.084 Quadrupole
- -1.2828 0.05 0.097 Quadrupole
- -0.876 0.4288 $w_{c}$ Plasmon
N=5 -1.067 0.882 0.0 G.S.
- -0.9377 0.0934 0.129 Quadrupole
- -0.562 0.396 $0.504\approx\omega_{c}$ Plasmon
- -0.091 0.086 0.543 Blue-shifted
@@
Conclusions.— Through exact-diagonalization, the laser-induced dynamics of FQH state at $\nu=1/3$ has been presented. We showed that laser-induced excitations of FQH states have rich physics. On the one hand, some excitations are coupled to electromagnetic field and can cause emission like the plasmon mode. On the other hand, we have shown that when scattering from LLL to LL1 is included in interactions, it is possible to excite neutral-modes of LLL through linearly polarized laser pulse field. These excitations are analogous to emergent gravitons due to their quadrupole structure and is the excitation of intrinsic metric carried by Hall fluids. Future experiments are needed to reveal how gap structure for fractional Hall systems evolve with finite momentum. Such experiments will not only address the questions related to strongly-interacting electron systems, but can also shed light on the understanding of the analogues of fundamental particles in high energy physics realized in condensed matter systems.
@@
Acknowledgements.— A.K. thanks Avadh Saxena and Pouyan Ghaemi for their fruitful discussions. This work was carried out under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration (NNSA) under Contract No. 89233218CNA000001. It was supported by LANL LDRD Program and Quantum Science Center, a U.S. DOE Office of Science Quantum Information Science Research Center. It was supported in part by Center for Integrated nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resources.

### v3

@@ Abstract
Where $X_{j}=\frac{2\pi j}{L_{y}}\ell_{B}^{2}$ with $\ell_{B}=\sqrt{\hbar/eB}$ being the magnetic length. ${\cal N}_{n}=L_{y}\sqrt{\pi}2^{n}n!\ell_{B}$ is the normalization constant and $H_{n}$ is Hermite polynomial of order $n$. For each Landau level $n$, index $j$ runs from $0,1,...,N_{\phi}-1$, where $N_{\phi}=L_{x}L_{y}/(2\pi\ell_{B}^{2})$ is the number of flux quanta (Landau orbitals) passing through the system. Within this Landau orbital basis, we are able to write the system Hamiltonian as $\hat{H}=\hat{H}_{sp}+\hat{H}_{int}$, with
@@
Figure 1: (a) Torus with lengths $L_{x}$ and $L_{y}$. Magnetic field is pointing along $\hat{z}$. (b) A schematic of a two-level FQH system. Initially, the system is in fractional $\nu=1/3$ state occupying the LLL ($n=0$). This state is then laser-driven creating excitations to the first Landau level ($n=1$). The energy difference between the two levels is $\omega_{c}$. (c) and (d) are the two-electron scattering processes that do not conserve Landau level index. (c) describes two electrons transitioning from one Landau level to another. (d) a two-electron process where only one electron scatters to another Landau level.
[…]
@@
Dynamical observables. Our system is initially in FQH-$\frac{1}{3}$ state $|\Psi_{0}\rangle$ at time $t=0$. We then turn on the laser drive and let the system evolve through Schrodinger equation as $|\Psi(t)\rangle=\hat{\mathit{T}}e^{-i\int_{0}^{t}\hat{H}(t^{\prime})dt^{\prime}}|\Psi_{0}\rangle$, where $\hat{\mathit{T}}$ is the time ordering operator and $|\Psi(t)\rangle$ is our out-of-equilibrium state. The time-ordering is achieved by discretizing the time-dependent matrix exponential in small time steps. The dynamical response of the system can be understood with time-dependent structure factor ${\mathcal{S}}(\mathbf{q})=\langle\hat{\rho}_{\mathbf{q}}\hat{\rho}_{-\mathbf{q}}\rangle$ where $\rho_{\mathbf{q}}$ is the projected density (see SM  sup ).
Results. We consider level-spacing $\omega_{c}<E_{c}$, i.e. the level spacing is less than Coulomb energy scale. The ratio of these energies $\kappa\equiv\frac{e^{2}}{\hbar\omega_{c}\epsilon\ell_{B}}$ is magnetic field dependent ranging from $2.6/\sqrt{B[T]}$ for Ga-As to $22.5/\sqrt{B[T]}$ for Al-As justifying the our approximation for fractional $\nu=1/3$ Hall system usually observed at $B\sim 10$ Tesla Sodemann and MacDonald (2013).
@@
Figure 2: Color online. Structure factor in frequency domain ${\mathcal{S}}_{\mathbf{q}}(\omega)$ for different system sizes. The torus dimensions is taken $L_{x}=L_{y}=\sqrt{6\pi N}\ell_{B}$. The laser-frequency is $\omega_{0}=0.7\omega_{c}$. The intensity is set to $eA_{0}/\sqrt{2}m^{*}\ell_{B}=0.2E_{c}$ and $t_{d}=10/E_{c}$. The units of time is the inverse of $E_{c}$. The usual values of $E_{c}$ is around $14\;{\rm meV}$ Pinczuk et al. (1993) which gives 4747 femto-seconds as the unit of time. The structure factor spectrum is normalized such that $\int{\mathcal{S}}_{\mathbf{q}}(\omega)d\omega=1$.
@@
The spectrum of structure factor ${\mathcal{S}}_{\mathbf{q}}(\omega)$ is shown in Fig. 2. The peak at the level-spacing (cyclotron-energy) $\omega_{c}$ is the plasmon equivalent of FQH state identified as $q\to 0$ limit of magnetoplasmon mode (see SM sup for other quantities). Significantly, there emerge new modes located at the energy around $0.2\omega_{c}$ for different system sizes. To determine the nature of theses new low-energy modes in $S_{\mathbf{q}}(\omega)$, we evaluate the overlap $|\langle{\Psi(t)}|n\rangle|$ to analyze the weight of low-lying eigen-states $|n\rangle$ contained in the photo-induced excited state $|\Psi(t)\rangle$ as shown in Table 1. We have found that low-energy excitations are dominated by quadrupole FQH graviton states. These states are identified by matrix-element $I_{n}=|\langle\Psi_{0}|\hat{O}^{(2)}|n\rangle|$, where $\hat{O}^{(2)}$ is the operator with quadrupole structure and $|\Psi_{0}\rangle$ is our initial ground state before the application of laser pulse. The specific form of $\hat{O}^{(2)}$ is given as Liou et al. (2019)
@@
Figure 3: Color online. (a) $I_{n}=|\langle\Psi_{0}|\hat{O}^{(2)}|n\rangle|$ for $N=7$ for $L_{y}=\sqrt{2\pi 21}\ell_{B}$. Where $|\Psi_{0}\rangle$ is the FQH-1/3 state and $|n\rangle$ is the $n$-th excited state of Hamiltonian in Eq. (3). The inset of (a) gives the transition matrix element between FQH-1/3 state and the eigen-states of LLL. (b) Same as (a) but for $N=6$ and $L_{y}=\sqrt{2\pi 18}\ell_{B}$. The level-spacing is set to $0.5E_{c}$. Grey vertical line on plots is at energy $0.14E_{c}$.
@@
Table 1: The contribution of states in the photo-induced state $|\Psi(t)\rangle$ after the pulse for N=N=5, 6, 7 and 8. $\Delta E$ denotes the energy difference from the ground state. Overlap $|\langle n|\Psi(t)\rangle|$ gives the contribution of individual eigenstate $|n\rangle$ of the system Hamiltonian with energy $E_{n}$. $\omega_{c}=0.5E_{c}$.
@@
Size  [^1] The parameters of the laser pulse are set to $(eA_{0}/\sqrt{2}m^{*}\ell_{B},\omega_{0})=(0.2E_{c},0.7\omega_{c})$ with $t_{d}=10E_{c}$ and $L_{y}=\sqrt{2\pi N_{\phi}}\ell_{B}$. Similar results have been found for observed for near-TT limit. $E_{n}$ ($E_{c}$) Overlap $\Delta E$ $(E_{c})$ Nature
N=5 -1.067 0.883 0.000 G.S.
- -0.938 0.097 0.129 Quadrupole
- -0.852 0.01 0.215 Quadrupole
- -0.562 0.4 $0.504\approx\omega_{c}$ Plasmon
Total weight 0.097 Quadrupole
$N=6$ -1.381 0.850 0.000 G.S.
- -1.296 0.095 0.085 Quadrupole
- -1.283 0.041 0.098 Quadrupole
- -1.235 0.028 0.146 Quadrupole
- -1.210 0.034 0.171 Quadrupole
- -1.176 0.014 0.205 Quadrupole
- -0.876 0.429 $w_{c}$ Plasmon
Total weight 0.212 Quadrupole
N=7  [^2] Due to large Hilbert space, for $N=7$, we only have low-energy quadrupole states by which we identify graviton modes in the structure factor. For $N=8$, we have only have eigenvectors corresponding to first 100 lowest energy eigenstates that include few graviton states. It is expected that for $N=8$ States that give more than $0.01$ overlap with $|\psi(t)\rangle$ are considered. -1.702 0.827 0.000 G.S.
- -1.606 0.058 0.096 Quadrupole
- -1.587 0.065 0.115 Quadrupole
- -1.583 0.046 0.119 Quadrupole
- -1.562 0.013 0.140 Quadrupole
- -1.548 0.030 0.161 Quadrupole
- -1.530 0.024 0.172 Quadrupole
- -1.2 - $w_{c}$ Plasmon
Total weight 0.236 Quadrupole
N=8  [^2] Due to large Hilbert space, for $N=7$, we only have low-energy quadrupole states by which we identify graviton modes in the structure factor. For $N=8$, we have only have eigenvectors corresponding to first 100 lowest energy eigenstates that include few graviton states. It is expected that for $N=8$ States that give more than $0.01$ overlap with $|\psi(t)\rangle$ are considered. -2.035 0.797 0.000 G.S.
- -1.951 0.026 0.084 Quadrupole
- -1.926 0.049 0.109 Quadrupole
- -1.909 0.063 0.126 Quadrupole
- -1.896 0.013 0.139 Quadrupole
- -1.884 0.038 0.151 Quadrupole
- -1.873 0.019 0.162 Quadrupole
Total weight 0.208 Quadrupole
In Fig. 3, we present the values of transition matrix element $\langle n|\hat{O}^{(2)}|\Psi_{0}\rangle$ (un-normalized) between FQH 1/3 state and $n$-th excited state $|n\rangle$ for two cases i) When $|n\rangle$ is the eigen-state of the system Hamiltonian involving all scattering processes and ii) when $|n\rangle$ is the eigen-state of LLL Hamiltonian where all scattering processes are confined to LLL. The transition matrix element $I_{n}$ represents the system transition rate due to an oscillating metric tensor analogous to the passing of a gravitation wave and hence usually termed as graviton Yang (2016). Fig.  3 (a) and (b) shows the presence of such mode(s) for various system sizes. This mode occurs for both liquid limit and near-TT limit (See SM sup ). As seen from Fig. 3, the energy of graviton states $E_{grav}$, when all scattering processes are included, is slightly lower than the graviton states of a system when all interaction processes are confined to LLL. Furthermore, we follow Ref. Liou et al., 2019 to analyze the handedness of this graviton and find that it has the structure of $(q_{x}-iq_{y})^{2}$, suggesting its chiral nature (See SM  sup for the detail of analysis.). Previously, it has also been proposed  Kirmani et al. (2022); Liu et al. (2018) that this low-energy mode can be excited via geometric quench Liu et al. (2021). A variant of such graviton mode called chiral graviton mode has been observed in experiments recently for fractional Hall states at filling fractional $\frac{1}{3}$ and $\frac{2}{5}$ fillings Liang et al. (2024). Chiral graviton modes in FQH system has been previously excited by tuning the laser-frequency near resonance conditions between valence and conduction band in 2D quantum wells. Our work has demonstrated that chiral graviton modes can also be exited when the a short-time laser is tuned near Landau level gap which is easily accessible experimentally. Furthermore, the excitation of graviton mode can be observed along with magneto-plasmon mode when all scattering processes are considered as seen from the structure factor.
Figure 4: The structure factor $S_{\mathbf{q}}(\omega)$ (unnormalized) for $\mathbf{q}=(2\pi/L_{x},0)$, $\omega=0.7\omega_{c}$ and $N=6$ electron system. All parameters are set according to the Fig. (2). Without the two-particle chiral excitations (Fig. 1 (c)) the intra-$L$ mode is suppressed. Inset gives the strength of the main peak of the low-energy (graviton) and plasmon mode against the intensity of the laser-pulse. Quadratic dependence and linear dependence is evident for the low energy mode and plasmon mode respectively. The peak value of the plasmon mode is divided by 5 for better fit.
The appearent drop in the dynamical structure factor when normalized in Fig. 2 between $N=6$ and $N=7$ is due to the presence of more quadrupole excitations present in $N=7$ photo-excited state $|\Psi(t)\rangle$. As seen from Fig. 3 (a) and (b), there are more graviton states for $N=7$ consistent with Ref. Liou et al., 2019. These states are also identified in Table. 1. Most importantly, the total low-energy quadrupole overlap in the photo-induced state does not drop with the system size as seen from Table. 1. We have shown that the quadrupole excitations occur up-to $N=8$.
The two-particle process that excites the low-energy mode(s) through photo excitation, is given in Fig. 1 (c). This process describes inter-level two-electron transitions from one Landau level into another thus requiring two-photons. Moreover, this process has amplitude that is chiral in nature $\frac{(q_{x}-iq_{y})^{2}}{2\sqrt{q_{x}^{2}+q_{y}^{2}}}e^{-\frac{q^{2}}{2}}c_{1,j_{1}}^{\dagger}c_{1,j_{2}}^{\dagger}c_{0,j_{3}}c_{0,j_{4}}+\mathrm{H.c.}$ (See SM sup ). As seen in Fig. 4, when such (chiral) two-particle excitations are omitted, there is a significant drop in the low-energy mode present in the structure factor. The dependence of the low-energy and the plasmon peak strength with the laser intensity in the limit $A_{0}\to 0$ is also shown in inset of Fig. 4. It is evident that in the low intensity regime, the magnetoplasmon and the low-energy mode follow the linear and quadratic dependence of the field strength respectively. This is in accordance with the fact that graviton couple with two-particle processes while magnetoplasmon depends upon single-photon process. Quantifying such quadratic dependence gives an alternate avenue to measure the graviton modes through non-linear absorption experiments. One can then use a circularly polarized probe light to experimentally detect the chirality of the excited graviton modes if needed.
@@
Conclusions.— Through exact-diagonalization, the laser-induced dynamics of FQH state at $\nu=1/3$ has been presented. We showed that laser-induced excitations of FQH states have rich physics. On the one hand, some excitations are coupled to electromagnetic field and can cause emission like the plasmon mode. On the other hand, we have shown that when scattering from LLL to LL1 is included in interactions, it is possible to excite neutral-modes of LLL through linearly polarized laser pulse field. These excitations are analogous to emergent gravitons due to their quadrupole structure and is the excitation of intrinsic metric carried by Hall fluids. Moreover, we have shown that non-linear light absorption experiments are an alternate route to the detection of such modes being quadratically dependent upon the light’s intensity. Future non-linear absorption experiments are needed to probe such excitations. This can also help with understanding how the gap structure for fractional Hall systems evolve with finite momentum. Such experiments will not only address the questions related to strongly-interacting electron systems, but can also shed light on the understanding of the analogues of fundamental particles in high energy physics realized in condensed matter systems.
@@
Acknowledgements.— This work was carried out under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration (NNSA) under Contract No. 89233218CNA000001. It was supported by Quantum Science Center, a U.S. DOE Office of Science Quantum Information Science Research Center, and in part by Center for Integrated nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resources. A.K. thanks Avadh Saxena and Pouyan Ghaemi for their fruitful discussions. A.K. was supported LANL LDRD Program.